# Efficient Algorithm Level Error Detection for Number-Theoretic Transform Assessed on FPGAs

### v1

@@ Abstract
Polynomial multiplication stands out as a highly demanding arithmetic process in the development of post-quantum cryptosystems. The importance of number-theoretic transform (NTT) extends beyond post-quantum cryptosystems, proving valuable in enhancing existing security protocols such as digital signature schemes and hash functions. Due to the potential for errors to significantly disrupt the operation of secure, cryptographically-protected systems, compromising data integrity, and safeguarding against side-channel attacks initiated through faults it is essential to incorporate mitigating error detection schemes. This paper introduces algorithm level fault detection schemes in NTT multiplication, representing a significant enhancement compared to previous research. We evaluate this through the simulation of a fault model, ensuring that the conducted assessments accurately mirror the obtained results. Consequently, we attain a notably comprehensive coverage of errors. Finally, we assess the performance of our efficient error detection scheme on FPGAs to showcase its implementation and resource requirements. Through implementation of our error detection approach on Xilinx/AMD Zynq Ultrascale+ and Artix-7, we achieve a comparable throughput with just a 9% increase in area and 13% increase in latency compared to the original hardware implementations.
Index Terms:
@@
Fault detection, fast Fourier transform (FFT), FPGA, number-theoretic transform (NTT).
I Introduction
@@
Fast Fourier Transform (FFT) [1] algorithms which are used to compute the Discrete Fourier Transform (DFT) have various applications, spanning from digital signal processing to the efficient multiplication of large integers. When the coefficients of the polynomial are specifically chosen from a finite field, the resulting transform is known as the Number Theoretic Transform (NTT) [2], and it can be computed using FFT algorithms designed for operations within this particular finite field. An efficient approach for polynomial multiplication, NTT holds significant importance in post-quantum cryptosystems, e.g., lattice-based cryptosystems which is regarded as a leading contender for quantum-secure public key cryptography, primarily because of its broad applicability and security proofs grounded in the worst-case hardness of established lattice problems. The NTT has proven to be a potent tool that facilitates the computation of this operation with quasi-polynomial complexity $O(n.lgn)$. Several recent studies have been conducted on optimizing NTT [3], [4], [5], [6], [7], [8], and [9]. In addition to polynomial multiplication, the use of NTT has the potential to significantly enhance existing schemes by improving their security parameters. NTT is widely used in signature schemes, hash functions, and identification schemes. As a result, incorporating efficient error detection mechanisms in NTT for polynomial multiplication will not only improve the security and reliability but also mitigate the risk of fault attacks in the respective algorithms in post-quantum cryptography. Moreover, such schemes could at least alleviate the attack surface or be add-ons to other approaches.
I-A Related works
We catogorize relevant studies into three subsets. The initial category centers on fault attacks and countermeasures applied to post-quantum cryptographic schemes, particularly those extensively utilizing NTT. Ravi et al. [10] presented the first fault injection analysis of the NTT. The authors found a significant flaw in the way NTT is implemented in the pqm4 library [11]. The identified vulnerability, referred to as twiddle-pointer, is exploited to demonstrate practical and effective attacks on Kyber and Dilithium [12]. Several other authors have reported employing fault injection on structured lattice-based schemes as a foundation for attacks [13], [14], [15], and [16] threatening the security of implementations.
@@ I-A Related works
The focus of the second category lies in offering fault detection methodologies pertaining to the algorithmic level of classical cryptographic schemes. In [17], the authors suggested an effective algorithmic-level error detection for the ECSM window method, aiming to identify both permanent and transient errors. In [18], the authors presented error detection schemes designed for the double-and-add-always and Montgomery ladder ECSM algorithm utilized in non-supersingular elliptic curves.
@@
The third category focuses on developing fault detection schemes specifically for NTT. In [19], the authors integrate the advantages of bitslicing, a software implementation technique where a datapath of an $n$-bit processor is treated as $n$ parallel single-bit datapaths, to devise a fault countermeasure for the NTT used in Dilithium [20]. Sarker et al. [21] and [22], presented error detection architectures of NTT based on recomputation with encoded operands. The authors achieved high error coverage with low area overhead. Despite the recomputation mechanism, the latency is elevated. We note that despite high error coverage, for fast implementations, these works might increase the total time to levels not acceptable.
I-B Our Major Contributions
@@
  • We have proposed an algorithm level fault detection scheme of NTT multiplication which is widely used in lattice based cryptography. We achieved high error coverage with less latency compared to previous works. This was achieved through the implementation of algorithm-level error detection for NTT section and partial recomputation for pre-computation sub-block. We performed simulations for single and burst fault injection using our proposed scheme. The simulation outcomes demonstrated that our approach is capable of detecting diverse types of faults with high error coverage and offers protection against the NTT fault attack presented in [10] and those with respective fault models.
@@
  • Our error detection method is deployed on Xilinx/AMD Zynq Ultrascale+ and Artix-7. The results of our implementations indicate that we can attain a significantly high level of error coverage with only a 13% increase in latency and a 9% area bloat.
II Preliminaries
II-A Number Theoretic Transform (NTT)
@@
The NTT of a sequence $a$ is derived as:
$\widetilde{a}=NTT(a)=\left[\begin{array}[]{cccc}\omega^{0}&\omega^{0}&...&\omega^{0}\\ \omega^{0}&\omega^{1}&...&\omega^{n-1}\\ \omega^{0}&\omega^{2}&...&\omega^{2(n-1)}\\ .&.&...&.\\ \omega^{0}&\omega^{n-1}&...&\omega^{(n-1)^{2}}\end{array}\right]*\left[\begin{array}[]{c}a(0)\\ a(1)\\ a(2)\\ .\\ a(n-1)\end{array}\right]$ (1)
@@
Algorithm 1 Iterative NTT [24]
Input: $a\in\mathbb{Z}_{q}$ of length $n=2^{k}$, $\omega$ is primitive $n$-th root of unity
Output: $\widetilde{a}=\text{NTT}(a)$ in bit-reversed order
1: $\widetilde{a}=a$
@@
7:                    $u=\widetilde{a}[k.m+j]$
8:                    $t\equiv\omega\,.\,\widetilde{a}[k.m+j+m/2]$ mod $q$
9:                    $\widetilde{a}[k.m+j]\equiv(u+t)$ mod $q$
10:                  $\widetilde{a}[k.m+j+m/2]\equiv(u-t)$ mod $q$
11: return $\widetilde{a}$
II-B Negative Wrapped Convolution
@@
Output: $\hat{a}=\text{pre-process}(a)$
1: $\hat{a}=a$
2: for $i=0$ to $n-1$:
3:       $\hat{a}[i]=a[i]\,.\,$$\psi^{i}$
4: return $\hat{a}$
III Proposed Error Detection Scheme
In this section, we present our error detection schemes on polynomial multiplication, $\text{NTT}(\hat{a})\,\Circle\,\text{NTT}(\hat{b})$ where $\hat{a}=\text{pre-process}(a)$ and $\hat{b}=\text{pre-process}(b)$. We have devised an error detection scheme at the algorithm level on the component-wise NTT multiplication sub-block, $\text{NTT}(\hat{a})\,\Circle\,\text{NTT}(\hat{b})$, as well as an error detection scheme for the pre-process function, $\hat{a}=\text{pre-process}(a)$, involving careful utilization of recomputation using shifted operands.
Figure 1: Concurrent error detection scheme for NTT operation which is used in [25].
III-1 Component-wise NTT Multiplication
Our focus is on $\text{NTT}(\hat{a})\,\Circle\,\text{NTT}(\hat{b})$ operation in this section. Jou et al. [25], proposed an algorithm level error detection scheme on FFT network. Their proposed error detection scheme is depicted in Fig. 1. In order to achieve an efficient scheme, we have adapted the aforementioned approach to component-wise NTT multiplication. The primary focus is to use an error detection scheme with small overhead and high error detection ratio with respect to efficient realizations of the NTT. From NTT definition, we can use the relationships presented below:
$\widetilde{a}(0)=\sum_{j=0}^{n-1}a(j)\omega^{0}=\sum_{j=0}^{n-1}a(j),\quad j=0,1,...,n-1$ (2)
$\widetilde{b}(0)=\sum_{j=0}^{n-1}b(j)\omega^{0}=\sum_{j=0}^{n-1}b(j),\quad j=0,1,...,n-1$ (3)

### v2

# Efficient Algorithm Level Error Detection for Number-Theoretic Transform used for Kyber Assessed on FPGAs and ARM

@@ Abstract.
Polynomial multiplication stands out as a highly demanding arithmetic process in the development of post-quantum cryptosystems. The importance of the number-theoretic transform (NTT) extends beyond post-quantum cryptosystems, proving valuable in enhancing existing security protocols such as digital signature schemes and hash functions. CRYSTALS-KYBER stands out as the sole public key encryption (PKE) algorithm chosen by the National Institute of Standards and Technology (NIST) in its third round selection, making it highly regarded as a leading post-quantum cryptography (PQC) solution. Due to the potential for errors to significantly disrupt the operation of secure, cryptographically-protected systems, compromising data integrity, and safeguarding against side-channel attacks initiated through faults it is essential to incorporate mitigating error detection schemes. This paper introduces algorithm level fault detection schemes in the NTT multiplication using Negative Wrapped Convolution and the NTT tailored for Kyber Round 3, representing a significant enhancement compared to previous research. We evaluate this through the simulation of a fault model, ensuring that the conducted assessments accurately mirror the obtained results. Consequently, we attain a notably comprehensive coverage of errors. Furthermore, we assess the performance of our efficient error detection scheme for Negative Wrapped Convolution on FPGAs to showcase its implementation and resource requirements. Through implementation of our error detection approach on Xilinx/AMD Zynq Ultrascale+ and Artix-7, we achieve a comparable throughput with just a 9% increase in area and 13% increase in latency compared to the original hardware implementations. Finally, we attained an error detection ratio of nearly 100% for the NTT operation in Kyber Round 3, with a clock cycle overhead of 16% on the Cortex-A72 processor.
@@
Fault detection, Fast Fourier Transform (FFT), Kyber, Negative Wrapped Convolution, Number-theoretic transform (NTT).
ccs: Security and privacy  Embedded systems security
1. Introduction
@@
Fast Fourier Transform (FFT) (Cochran et al., 1967) algorithms which are used to compute the Discrete Fourier Transform (DFT) have various applications, ranging from digital signal processing to the efficient multiplication of large integers. When the coefficients of the polynomial are specifically chosen from a finite field, the resulting transform is known as the Number Theoretic Transform (NTT) (Cooley and Tukey, 1965), and it can be computed using FFT algorithms designed for operations within this particular finite field. An efficient approach for polynomial multiplication, NTT holds significant importance in post-quantum cryptosystems, e.g., lattice-based cryptosystems which are regarded as a leading contender for quantum-secure public key cryptography, primarily because of its broad applicability and security proofs grounded in the worst-case hardness of established lattice problems.
In 2022, NIST unveiled the Round 3 outcomes of the Post-Quantum Cryptography standardization Process, which featured four chosen algorithms (CRYSTALS-KYBER , CRYSTALS-DILITHIUM, Falcon, SPHINCS+). These are called formally after Aug. 2024 with respect to Federal Information Processing Standards (FIPS) as FIPS 203, Module-Lattice-Based Key-Encapsulation Mechanism Standard, FIPS 204, Module-Lattice-Based Digital Signature Standard, and FIPS 205, Stateless Hash-Based Digital Signature Standard while Falcon is yet to get a FIPS number (NIST, 2024a), (NIST, 2024b), and (NIST, 2024c). Out of the chosen algorithms, CRYSTALS-KYBER, also referred to as Kyber (Bos et al., 2018) and (NIST, 2024a), stands out as the sole algorithm for both public key encryption and key establishment that relies on the challenge of the module learning with errors (MLWE) problem.
The NTT has proven to be a potent tool that facilitates the computation of this operation with quasi-linear complexity $O(n.lgn)$. Several recent studies have been conducted on optimizing the NTT (Bisheh-Niasar et al., 2021), (Geng et al., 2023),(Hermelink et al., 2023), (Hwang et al., 2022), (Kurniawan et al., 2023), (Li et al., 2023), (Mu et al., 2023), (Tosun and Savas, 2024), and (Zhang et al., 2021). In addition to polynomial multiplication, the use of the NTT has the potential to significantly enhance existing schemes by improving their security parameters. NTT is widely used in signature schemes, hash functions, and identification schemes. As a result, incorporating efficient error detection mechanisms in the NTT for polynomial multiplication will not only improve the security and reliability but also mitigate the risk of fault attacks in the respective algorithms in post-quantum cryptography. Moreover, such schemes could at least alleviate the attack surface or be add-ons to other approaches. Kyber and Dilithium employ polynomial multiplication over $\mathbb{Z}[X]/(X^{n}+1)$ utilizing the NTT. Integrating effective error detection schemes into the NTT used in Kyber will enhance security, reliability, and mitigate the risk of fault attacks in post-quantum cryptography algorithms.
1.1. Related works
Several previous studies have concentrated on the implementation and fault detection in different arithmetic components of both classical and post-quantum cryptography (Canto et al., 2022a), (Guo et al., 2024), (Libano et al., 2023), (Richter-Brockmann et al., 2020), and (Sarker et al., 2022b). We categorize relevant studies into three subsets. The initial category centers on fault attacks and countermeasures applied to post-quantum cryptographic schemes, particularly those extensively utilizing the NTT. Numerous prior studies have explored fault attacks with the objective of gaining access to or compromising secret keys (ElGhamrawy et al., 2023), (Heinz and Pöppelmann, 2022), (Hermelink et al., 2021), (Ravi et al., 2024), (Saha et al., 2023), and (Xagawa et al., 2021). Ravi et al. (Ravi et al., 2023) presented the first fault injection analysis of the NTT. The authors found a significant flaw in the way the NTT is implemented in the pqm4 library (Kannwischer et al., 2019). The identified vulnerability, referred to as twiddle-pointer, is exploited to demonstrate practical and effective attacks on Kyber and Dilithium (Avanzi et al., 2019). In (Ravi et al., 2019), the authors proposed a fault attack aiming the NTT operation during the key generation and encapsulation routine of the Kyber. The fault attack is achieved by zeroize all the twiddle factors used in the NTT operation. When focusing on applying this technique to the NTT regarding secrets or errors, it can significantly diminish the randomness of the secret or error and reduce the entropy. Several other authors have reported employing fault injection on structured lattice-based schemes as a foundation for attacks (Bindel et al., 2016), (Bruinderink and Pessl, 2018), and (Mus et al., 2020) threatening the security of implementations. Regarding active attacks, although some prior research addressed fault injection, Espitau et al. (Espitau et al., 2018) introduced loop-abort faults in several lattice-based cryptosystems, including CRYSTALS-Kyber. In this attack, a fault is injected into the cryptosystem, causing the loop responsible for sampling random Gaussian secret coefficients to terminate early. This early termination leads to the generation of lower-dimensional secrets, which can then be leveraged for a key recovery attack. In 2021, Pessl and Prokop (Prokop and Peßl, 2021) proposed an attack that involves a single instruction-skipping fault during the decoding process. Their fault simulations showed that at least 6,500 faulty decapsulation attempts are needed to fully recover the key for Kyber512 operating on a Cortex M4. In the same year, Hermelink et al. (Hermelink et al., 2021) combined fault injections with chosen-ciphertext attacks on CRYSTALS-Kyber, suggesting that their attack could bypass defenses such as decoder shuffling. Their findings demonstrated successful secret key recovery using 7,500 inequalities for Kyber-512, 10,500 for Kyber-768, and 11,000 for Kyber-1024. Delvaux (Delvaux, 2021) refined the side-channel attack (SCA) from (Hermelink et al., 2021), making it simpler to execute and more challenging to defend against. In (Kundu et al., 2024), the authors proposed a new fault attack on the SCA-secure masked decapsulation algorithm for generic LWE-based KEMs and detailed the attack for Kyber. Jendral (Jendral, 2024) presented an attack on CRYSTALS-Dilithium implementation on ARM Cortex-M4 using fault injection. Krahmer et al. (Krahmer et al., 2024) presented two key-recovery fault attacks on Dilithium’s signing procedure.
@@ 1.1. Related works
The focus of the second category lies in offering fault detection methodologies pertaining to the algorithmic level of classical cryptographic schemes. In (Ahmadi et al., 2024b), the authors suggested an effective algorithmic-level error detection for the ECSM window method, aiming to identify both permanent and transient errors. In (Ahmadi et al., 2024a), the authors presented algorithm level error detection scheme designed for Montgomery ladder ECSM algorithm utilized in non-supersingular elliptic curves.
@@
The third category focuses on developing fault detection schemes specifically for the NTT. In (Bauer et al., 2024), the authors introduced a method for safeguarding the NTT against fault attacks. Mishra et al. (Mishra and Mukhopadhyay, 2024) introduced a principle for countermeasures that is based on cryptographic guarantees rather than relying on ad hoc methods, aiming to offer measurable protection against the previously mentioned fault attacks on lattice-based schemes. In (Singh et al., 2024), the authors integrate the advantages of bit slicing, a software implementation technique where a datapath of an $n$-bit processor is treated as $n$ parallel single-bit datapaths, to devise a fault countermeasure for the NTT used in Dilithium (Ducas et al., 2018). Sarker et al. (Sarker et al., 2022a) and (Sarker et al., 2018), presented error detection architectures of the NTT based on recomputation with encoded operands. The authors achieved high error coverage with low area overhead. However, due to the recomputation process, their latency is doubled. We note that despite high error coverage, for fast implementations, these works might increase the total time to levels not acceptable. Additionally, Cintas-Canto et al. (Canto et al., 2022b) introduced error detection schemes for lattice-based KEMs using recomputation techniques and implemented these schemes on FPGA. Below, we present our major contributions in this work.
1.2. Our Major Contributions
  • Our proposed error detection schemes donot depend on recomputation; instead, it relies on algorithm-level coherency. As a result, it offers higher speed, lower latency, and reduced area compared to previous approaches.
@@ 1.2. Our Major Contributions
  • We have proposed an algorithm level fault detection scheme of the NTT multiplication using Negative Wrapped Convolution which is widely used in lattice-based cryptography. We achieved high error coverage with less area overhead and latency compared to previous works. This was achieved through the implementation of algorithm-level error detection for the NTT section and partial recomputation for pre-computation section.
  • As the Negative Wrapped Convolution method does not work for the NTT multiplication used in Kyber Round 3, we have introduced an algorithm level error detection scheme for the NTT multiplication tailored for Kyber Round 3. Implementing algorithm level error detection for the NTT module enabled us to achieve substantial error coverage with minimal overhead compared to previous works. This method can be integrated into Kyber Round 3 reference implementation.
  • We performed simulations for single and burst fault injection using our proposed schemes. The simulation outcomes demonstrated that our approach is capable of detecting diverse types of faults with high error coverage and offers protection against the NTT fault attack presented in (Ravi et al., 2023) and those with respective fault models.
@@
  • Our error detection method for Negative Wrapped Convolution was deployed on Xilinx/AMD Zynq Ultrascale+ , and Artix-7. The results of our implementations indicate that we can attain a significantly high level of error coverage with only a 13% increase in latency and a 9% area bloat. Our proposed error detection method for Kyber Round 3 is implemented on Intel Core-i7, and Cortex-A72 with additional overhead of 28% and 16% in terms of clock cycle overhead.
Our work is structured as follows: Section 2 discusses Negative Wrapped Convolution and the NTT used in Kyber. The presented error detection schemes on Negative Wrapped Convolution and the NTT tailored for Kyber are described in Section 3. Section 4 is divided into two subsections. First, the fault model in this work is discussed; next, the fault simulation is performed to assess error detection rate. We apply such fault detection schemes into the reference Kyber implementation and Negative Wrapped Convolution in Section 5 to benchmark the different overheads. Finally, our conclusions are presented in Section 6.
2. Preliminaries
2.1. Number Theoretic Transform (NTT)
@@
(1) $\hat{f}=\text{NTT(}f)=\left[\begin{array}[]{cccc}\omega^{0}&\omega^{0}&...&\omega^{0}\\ \omega^{0}&\omega^{1}&...&\omega^{n-1}\\ \omega^{0}&\omega^{2}&...&\omega^{2(n-1)}\\ .&.&...&.\\ \omega^{0}&\omega^{n-1}&...&\omega^{(n-1)^{2}}\end{array}\right]*\left[\begin{array}[]{c}f(0)\\ f(1)\\ f(2)\\ .\\ f(n-1)\end{array}\right].$
@@
Algorithm 1 Iterative NTT (Zhang et al., 2019)
Input: $f\in\mathbb{Z}_{q}$ of length $n=2^{k}$, $\omega$ is primitive $n$-th root of unity
Output: $\hat{f}=\text{NTT}(f)$ in bit-reversed order
1: $\hat{f}=f$
@@
7:                    $u=\hat{f}[k.m+j]$
8:                    $t\equiv\omega\,.\,\hat{f}[k.m+j+m/2]$ mod $q$
9:                    $\hat{f}[k.m+j]\equiv(u+t)$ mod $q$
10:                  $\hat{f}[k.m+j+m/2]\equiv(u-t)$ mod $q$
11: return $\hat{f}$
2.2. Negative Wrapped Convolution
@@
Algorithm 2 The pre-process step for the NTT calculations
[…]
Output: $\tilde{f}=\text{pre-process}(f)$
1: for $i=0$ to $n-1$:
2:       $\tilde{f}[i]=f[i]\,.\,$$\psi^{i}$
3: return $\tilde{f}$
Algorithm 3 The Post-process step for the NTT calculations
[…]
Output: $\tilde{f}=\text{post-process}(f)$
1: for $i=0$ to $n-1$:
2:       $\tilde{f}[i]=f[i]\,.\,$$\psi^{-i}$
3: return $\tilde{f}$
Algorithm 4 Kyber NTT
Input: $r\in\mathbb{Z}_{q}$ of length $n=256$, $\omega=17$ is primitive $n$-th root of unity
Output: $\text{NTT}(r)$ in bit-reversed order
1: $\widetilde{r}=r$
2: $j=0$
3: $k=0$
4: for ($s=128;s>=2;s=s/2$):
5:       for ($i=0;i<256;i=j+s$):
6:              $zeta=\omega^{\text{(bit-reverse}(k))}$
7:              $k=k+1$
8:              for $(j=i;j<i+s;j++)$
9:                    $t=\text{fqmul}(zeta,\widetilde{r}[j+s])$
10:                   $\widetilde{r}[j+s]=\widetilde{r}[j]-t$ mod $q$
11:                  $\widetilde{r}[j]=\widetilde{r}[j]+t$ mod $q$
12: return $\widetilde{r}$
2.3. Polynomial Multiplication Utilizing the NTT in Kyber
To use Negative Wrapped Convolution method, we require that $2n\mid(q-1)$. Regarding the parameters used in Kyber with the prime $q=3329$ and $n=256$ the base field $\mathbb{Z}_{q}$ contains 256-th roots of unity but not 512-th roots. Therefore, we cannot use the method used in (Lyubashevsky et al., 2008). To overcome this problem, the polynomial $X^{256}+1$ of $R$ factor into 128 polynomials of degree 2 modulo $q$. The polynomial $X^{256}+1$ can be written as
$X^{256}+1=\prod_{i=0}^{127}(X^{2}-\omega^{2i+1})=\prod_{i=0}^{127}(X^{2}-\omega^{2b_{7}(i)+1}),$
where $b_{7}(i)$ is the bit reversal of the 7-bit $i$. Therefore, the NTT of $f$ is a vector of 128 polynomials of degree one.
$\hat{f}=\text{NTT(}f)=(\hat{f}_{0}+\hat{f}_{1}X,\hat{f}_{2}+\hat{f}_{3}X,...,\hat{f}_{254}+\hat{f}_{255}X),$
with
(2) $\hat{f}_{2i}=\sum_{j=0}^{127}f_{2j}\omega^{(2br_{7}(i)+1)j},$
(3) $\hat{f}_{2i+1}=\sum_{j=0}^{127}f_{2j+1}\omega^{(2br_{7}(i)+1)j}.$
[…]
(4) $\hat{h}_{2i}+\hat{h}_{2i+1}X$ $=(\hat{f}_{2i}+\hat{f}_{2i+1}X)(\hat{g}_{2i}+\hat{g}_{2i+1}X)\text{ mod ($X^{2}$-$\omega^{2b_{7}(i)+1}$)}.$
The iterative NTT implementation of Kyber is described in Algorithm 4. fqmul function (presented in Line 9 of Algorithm 4) performs multiplication of two operand and Montgomery reduction afterward.
3. Proposed Error Detection Scheme
3.1. Negative Wrapped Convolution
In this section, we present our error detection schemes on the sub-block in the Negative Wrapped Convolution that uses the NTT and pre-process modules, $\text{NTT}(\tilde{f})\,\Circle\,\text{NTT}(\tilde{g})$ where $\tilde{f}=\text{pre-process}(f)$ and $\tilde{g}=\text{pre-process}(g)$. We have devised an error detection scheme at the algorithm level on the component-wise NTT multiplication sub-block, $\text{NTT}(\tilde{f})\,\Circle\,\text{NTT}(\tilde{g})$, as well as an error detection scheme for the pre-process function, involving careful utilization of recomputation using shifted operands. This error detection scheme can be applied to designs which require polynomial multiplication using the NTT. We used Negative Wrapped Convolution in our proposed error detection scheme.
3.1.1. Component-wise NTT Multiplication
Our focus is on $\text{NTT}(\tilde{f})\,\Circle\,\text{NTT}(\tilde{g})$ operation in this section. Jou et al. (Jou and Abraham, 1988) proposed an algorithm level error detection scheme on FFT network. Their proposed error detection scheme is depicted in Fig. 1. In order to achieve an efficient scheme, we have adapted the aforementioned approach to the component-wise NTT multiplication. The primary focus is to use an error detection scheme with small overhead and high error detection ratio with respect to efficient realizations of the NTT. From the NTT definition, we can use the relations presented below:
Figure 1. Concurrent error detection scheme for the NTT operation.
(5) $\hat{f}(0)=\sum_{j=0}^{n-1}\tilde{f}(j)\omega^{0}=\sum_{j=0}^{n-1}\tilde{f}(j),\quad j=0,1,...,n-1,$
(6) $\hat{g}(0)=\sum_{j=0}^{n-1}\tilde{g}(j)\omega^{0}=\sum_{j=0}^{n-1}\tilde{g}(j),\quad j=0,1,...,n-1.$